OpenStudy (anonymous):

What is the inverse of this function? f(x)=Log2(7x) ((2 is the base of log))

OpenStudy (anonymous):

y=2^x/7

OpenStudy (anonymous):

(2^x)/7 There

OpenStudy (anonymous):

you can check too: f(f-1(x))=Log2(7*(2^x)/7) =Log2(2^x) =x

OpenStudy (anonymous):

start by switching the x & y's so it turns to x=log2(7y). then use the rule to make 2^x =7y. then divide by 7.